# Deblending the MIGHTEE-COSMOS survey with XID+: The resolved radio source counts to Jy

### v1

@@ Appendix A Corner Plots for Star-Galaxy Classification
To supplement the two-dimensional UMAP projections shown in the main text, Figures 18 shows the corner plot for all the 10-dimensional reduced feature space for the HSC+VISTA dataset. This plot illustrate how the separation between the stellar and galactic clusters is maintained across the various pairings of the reduced UMAP dimensions, reinforcing the robustness of the classification.
@@ Appendix B Deblending Performance with GAUL-based Masking
To supplement the primary analysis presented in Section 4.4.4, we also investigated the impact of using a bright-source mask generated from the PyBDSF GAUL catalog instead of the SRL (source list) catalog. The GAUL catalog treats every detected Gaussian component individually, making it a useful alternative to test the robustness of our masking strategy.
The deblending performance plots for this test are presented in Figures 19 and 20. The results are qualitatively and quantitatively very similar to those derived using the SRL mask in the main text. This confirms that our main conclusion, that the quality of the prior list and the 50 Jy masking threshold are the dominant factors driving the fidelity of the deblended flux densities is robust and not sensitive to the specific catalog type used to generate the bright-source mask.
@@ Appendix C Deblending with a Comprehensive LIR Prior
To provide a real-world validation of our simulation results (Sec. 4), we also performed a deblending run on the MIGHTEE data using Comprehensive LIR Prior. This prior was designed to be as inclusive as possible, containing 298,894 potential sources after masking, and is analogous to the "Full Prior" used in our simulations. Using the same optimized sampler settings as for the Radio-Likely Prior (max_treedepth up to 15), we find that 6,675 sources have a poor fit (p-value > 0.6), a dramatic reduction from the 22,133 flagged in the initial run. However, even with optimized settings, a prominent cloud of sources with significantly overestimated flux densities remains when compared to the independent super-deblended catalog. While the fraction of flagged sources has dropped to 2.2 per cent of the total sample, the overall scatter in the flux comparison is visibly larger than in the Radio-Likely Prior case.
@@ Appendix D Supplementary Data Tables
The data tables for the derived source counts and a description of the final deblended source catalog.
@@ D.1 Radio Source Counts Data
The Euclidean-normalized source counts derived from the final, quality-filtered Radio-Likely Prior catalog are provided in Table 4. These data points are plotted in Figure 17.
@@
D.2 Deblended Catalog Format
The format and column descriptions for the final deblended MIGHTEE radio point source catalog are provided in Table 5.

### v2

@@ Appendix A Corner Plots for Star-Galaxy Classification
To supplement the two-dimensional UMAP projections shown in the main text, Figure 18 shows the corner plot for all the 10-dimensional reduced feature space for the HSC+VISTA dataset. This plot illustrate how the separation between the stellar and galactic clusters is maintained across the various pairings of the reduced UMAP dimensions, reinforcing the robustness of the classification.
@@ Appendix B Deblending Performance with GAUL-based Masking
To supplement the primary analysis presented in Section 4.4.4, we also investigated the impact of using a bright-source mask generated from the PyBDSF GAUL catalogue instead of the SRL (source list) catalogue. The GAUL catalogue treats every detected Gaussian component individually, making it a useful alternative to test the robustness of our masking strategy.
The deblending performance plots for this test are presented in Figures 19 and 20. The results are qualitatively and quantitatively very similar to those derived using the SRL mask in the main text. This confirms that our main conclusion, that the quality of the prior list and the 50 Jy masking threshold are the dominant factors driving the fidelity of the deblended flux densities is robust and not sensitive to the specific catalogue type used to generate the bright-source mask.
@@ Appendix C Deblending with a Comprehensive LIR Prior
To provide a real-world validation of our simulation results (Sec. 4), we also performed a deblending run on the MIGHTEE data using Comprehensive LIR Prior. This prior was designed to be as inclusive as possible, containing 298,894 potential sources after masking, and is analogous to the Full Prior used in our simulations. Using the same optimized sampler settings as for the Radio-Likely Prior (max_treedepth up to 15), we find that 6,675 sources have a poor fit (p-value > 0.6), a dramatic reduction from the 22,133 flagged in the initial run. However, even with optimized settings, a prominent cloud of sources with significantly overestimated flux densities remains when compared to the independent super-deblended catalogue. While the fraction of flagged sources has dropped to 2.2 per cent of the total sample, the overall scatter in the flux comparison is visibly larger than in the Radio-Likely Prior case.
@@ Appendix D Supplementary Data Tables
The data tables for the derived source counts and a description of the final deblended source catalogue.
@@ D.1 Radio Source Counts Data
The Euclidean-normalized source counts derived from the final, quality-filtered Radio-Likely Prior catalogue are provided in Table 4. These data points are plotted in Figure 17.
@@
D.2 Deblended Catalogue Format
The format and column descriptions for the final deblended MIGHTEE radio point source catalogue are provided in Table 5.